whats the x and y intercepts for -x+6y=20
Answer by jim_thompson5910(35256)

x-intercept
-x+6y=20

-x+6(0)=20

-x=20

x=-20

So the x-intercept is (-20, 0)

---------------------------------------------
y-intercept


-x+6y=20

-(0)+6y=20

6y=20

y = 20/6

y = 10/3

y = 3.3333

So the y-intercept is (0, 3.3333)